OpenStudy (anonymous):

2 is added to m and the square root of the sum is taken. Then, 7 is subtracted from the result, giving a final result of 9. What is m?

OpenStudy (cwrw238):

sqrt(m + 2) - 7 = 9

OpenStudy (anonymous):

\[\sqrt{m+2}-7=9\]

OpenStudy (anonymous):

\[\sqrt{m+2}=16\]